OpenStudy (anonymous):

Find the area common to the 2 regions bounded by: r = - 6 cos(x)  r = 2 - 2 cos(x) 

OpenStudy (anonymous):

First area of the cardioid \[ \text{area3}=\int_0^{2 \pi } \frac{1}{2} (2-2 \cos (x))^2 \, dx= 6\pi \]

OpenStudy (anonymous):

Then the area outside the circle but within the cardioid \[ 2 \left(\int_0^{\frac{2 \pi }{3}} \frac{1}{2} (2-2 \cos (x))^2 \, dx-\int_{\frac{2 \pi }{3}}^{\frac{\pi }{2}} \frac{36 \cos ^2(x)}{2} \, dx\right)=\\2\left ( 2 \pi -\frac{9 \sqrt{3}}{4} -\frac{3}{4} \left(3 \sqrt{3}-2 \pi \right) \right)=7 \pi -9 \sqrt{3}\\ \]

OpenStudy (anonymous):

The area in question is \[ 6\pi - 7 \pi + 9 \sqrt 3=9 \sqrt{3}-\pi \]